In a binomial distribution the probability of getting zero or more number of successes is equal to

a. 0

b. the probability of getting zero success

c. the probability of getting successes in all the trials

d. 1 minus the probability of getting successes in all the trials

e. 1


1
Expert's answer
2021-03-01T15:02:31-0500

Let the probability of getting success be p.

So the probability of getting fail q= 1 -p

So, probability of getting zero success "P(B)=q^n"

Probability of getting at least one success "P(A)=" "1-(q)^n"

So, the probability of getting zero or more number of success "=P(A)+P(B)"

"=(1-q^n)+q^n"

Here,the probability to get zero success and the probability to get at least one success are mutually exclusive events.

And we know that the sum of all the probability always equal to 1.

So, correct answer will be option e.